OpenStudy (anonymous):

In 5 years, Dad will be three times as old as his daughter Jill will be then. If the sum of their present ages is 50, how old are they now? If x + 5 is Jill's age five years from now, which of the following equations could be used to solve the problem? a.3x + 5 = 55 - x b.3x + 15 = 55 - x c.3x + 15 = 45 - x

OpenStudy (anonymous):

y+5=3(x+5) x+y=50

OpenStudy (anonymous):

y+5=3x+15 x+y=50 y=-x+50 (-x+50)+5=3x+15

OpenStudy (anonymous):

-x+55=3x+15